# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### 04 Jan 2024 v3

@@ V Results
V-A Performance of the Proposed CLIP-based Frameworks
We present the performance summary of various CLIP-based frameworks on the four distracted driving datasets in Table II. The VideoCLIP model with an optimal hyperparameter setting demonstrated the most superior performance with a Top-1 accuracy of 81.85%, 98.44% and 97.86% on SynDD1, DMD and SAM-DD datasets, respectively. This was followed by the Multi-frameCLIP framework with an optimal hyperparameter setting. Single-frameCLIP and Zero-shotCLIP reported lower Top-1 accuracies on the SynDD1 dataset compared to the Statefarm dataset.
TABLE II: Performance Summary of CLIP-based
@@ V-A Performance of the Proposed CLIP-based Frameworks
(ours) Top-1 Top-3 Top-1 Top-3 Top-1 Top-3 Top-1 Top-3
Zero-shotCLIP 54.00 74.50 58.22 67.63 48.12 79.56 66.52 86.33
Single-frameCLIP 57.35 82.25 83.15 95.75 82.65 97.76 89.14 98.78
Multi-frameCLIP 70.36 75.38 N/A N/A 93.39 95.74 89.14 90.27
VideoCLIP 81.85 91.96 N/A N/A 98.44 99.80 97.86 99.76
On the StateFarm dataset, only Single-frameCLIP and Zero-shotCLIP were evaluated due to the absence of video data, which is a prerequisite for Multi-frameCLIP and VideoCLIP. Single-frameCLIP outperformed Zero-shotCLIP with optimal hyperparameter setting, achieving a Top-1 accuracy of 83.15% and a Top-3 accuracy of 95.75%.
These results highlight the effectiveness of VideoCLIP and Multi-frameCLIP in leveraging temporal information from video data for improved performance. It is important to note that we conduct comprehensive testing across various hyperparameter settings and select the parameters that deliver the highest performance for our proposed models. We refer to this setting as the “optimal hyperparameter setting”. A detailed analysis of these results and a discussion of the optimal hyperparameter setting are presented in the following sections.
1)  Zero-shotCLIP results
For the zero-shot experiments, we test the Zero-shotCLIP model on the distracted driving dataset without any prior training. The model encounters previously unseen frames extracted from distracted driving datasets. The RGB input images are resized to a resolution of 224x224x3. During the zero-shot evaluation, the framework extracts visual and textual features from the input images and corresponding textual prompts. Subsequently, the cosine similarity between these visual and textual features is computed, resulting in a ranked list of the most probable classes based on their similarity scores. The zero-shot transfer results are presented in Table III for different CLIP backbones. It can be seen that the ViT-L/14@336px vision transformer-based Zero-shotCLIP outperforms other variants in terms of Top-1 accuracy. It should be noted that the ViT-L/14@336px backbone processes input images at a higher RGB resolution (336x336x3). A key observation from this table is that ViT-L/14@336px based Zero-shotCLIP reached up to 86.33% Top-3 accuracy in the SAM-DD dataset without requiring dataset-specific training.
TABLE III: Zero-shot transfer results
CLIP SynDD1 DMD SAM-DD
backbones Top-1 Top-3 Top-1 Top-3 Top-1 Top-3
ViT-L/14 53.50 74.50 47.97 79.56 62.51 84.52
ViT-L/14@336px 54.00 65.00 48.12 77.16 66.52 86.33
ViT-B/16 47.50 64.50 31.29 54.95 18.67 79.18
ViT-B/32 38.00 45.50 28.39 47.76 10.48 40.09
2)  Single-frameCLIP results
We demonstrate the performance of the Single-frameCLIP model with optimal hyperparameter setting on the four datasets in Table IV. Single-frameCLIP results showcase the advantage of task-based fine-tuning over Zero-shotCLIP. For the SAM-DD dataset, the Single-frameCLIP model with the ViT-L/14 backbone achieved a Top-1 accuracy of 89.14% with the side-view frames. On the SynDD1 dataset, the same model reached a Top-1 accuracy of 57.35%. The results further confirm the robustness and effectiveness of the ViT-L/14-based backbones in the Single-frameCLIP model for distracted driving recognition tasks. It should be noted that the ‘leaving a set of driver out’ and the $k$-fold cross-validation method played a vital role in determining the optimal hyperparameter setting.
TABLE IV: Single-frameCLIP results with Optimal Hyperparameters
Single-frameCLIP Dataset Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 StateFarm Side-view 83.15
ViT-L/14 SynDD1 Dashboard 57.35
ViT-L/14@336px DMD Side-view 82.65
ViT-L/14 SAM-DD Side-view 89.14
3)  Multi-frameCLIP results
The Multi-frameCLIP framework involves employing majority voting over a set of consecutive frames to determine the final prediction of an action in a video clip. We investigate various sampling rates of 0.5, 1, 2, 5, and 20 frames per second, subject-level cross-validation along with different camera views and CLIP backbones. The result with the optimal setting is presented in Table V. The best performance is achieved with the ViT-L/14@336px backbone on the DMD dataset for the side-view video clip sampled at 1 frame per second. This optimal setting provided a Top-1 accuracy of 93.39%. We have observed that the model achieves peak performance at a sampling rate of 1 FPS. Any increase in the sampling rate resulted in a decline in accuracy (as shown in Table XIII) since the inclusion of similar-looking frames acts as a form of noisy signal in the majority voting pool. On the other hand, reducing the sampling rate below 1 compromised the performance since critical temporal information that is essential for final action prediction gets overlooked. Therefore, we can conclude that 1 FPS is the optimal sampling rate for the Multi-frameCLIP model to adequately capture relevant temporal information. The main advantage of this model is that it processes several frames simultaneously, exploiting the temporal information between two consecutive frames. This approach, combined with the high-level semantic understanding of the CLIP model, resulted in an improved performance. The optimal hyperparameters were determined through a series of experiments to optimize the model’s overall performance.
TABLE V: Multi-frameCLIP results with Optimal Hyperparameters
Multi-frameCLIP Dataset FPS Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 SynDD1 1 Dashboard 70.36
ViT-L/14@336px DMD 1 Side-view 93.39
ViT-L/14 SAM-DD 1 Side-view 89.14
4)  VideoCLIP results
TABLE VI: VideoCLIP results with Optimal Hyperparameters
VideoCLIP Dataset FPS Camera Top-1 Leaving a set
backbone view accuracy of driver out
ViT-L/14 SynDD1 20 Dashboard 81.85
ViT-L/14@336px DMD 1 Side-view 98.44
ViT-L/14 SAM-DD 1 Side-view 97.86
Similar to Multi-frameCLIP, we conduct a series of tests to obtain the optimal hyperparameter setting for the VideoCLIP model and report the results in Table VI. The best performance of 97.86% Top-1 accuracy is achieved on the SAM-DD dataset with the ViT-L/14 backbone and with a sampling rate of 1 frame per second. The performance achieved by the VideoCLIP model validates its efficacy in understanding distracted actions from driving video footage. We have observed that the sampling rate has a significant effect on the final accuracy. Moreover, in contrast to Multi-frameCLIP, VideoCLIP uses a sliding window of 30 frames and employs a majority voting scheme on the outputs of these small video clips. It is important to note that on the SynDD1 dataset, the VideoCLIP achieved the best performance at 20 FPS sampling rate. If the sampling rate is too small, e.g., 1 FPS, there will be one frame sampled from every second of an activity. Since VideoCLIP uses 30 frames to build a video clip, this clip will span 30 seconds which can be longer than the average action duration (from Fig. 5 the average action duration for SynDD1 dataset is $\sim$18.7 sec) and confuse the network. For example, an action video lasting 20 seconds should contain 20 frames at 1FPS; however, in these 20 frames, only 10 frames contain valid information about that action. Now, if we sample 30 frames from those 20 frames to make up a window as the input to the VideoCLIP network, most frames in that window are noise. In addition, the sliding window approach in VideoCLIP provides a more holistic representation of the video data, which leads to improved prediction accuracy. In the following sections, we will delve into the performance of our models under real-world variabilities and scrutinize the model’s efficacy across diverse scenarios.
V-B Performance Comparison with Other Works
@@
In this section, we compare the performance of the Single-frameCLIP model with traditional distracted driving backbones on frame-based datasets. We provided two distinct tables to highlight how the accuracy scores change across different test driver sets and model backbones on the StateFarm and SynDD1 datasets. The Top-1 accuracy scores from our runs are presented in Tables VIIVIII and Fig. 6. It should be noted that, across all model comparisons in our experiments, we maintain consistency by using the same experimental settings along with the same action classes. Table VII shows the detailed breakdown of our ‘leave a set of drivers out’ approach that we used for testing on the StateFarm dataset. The first column entries represent the drivers included in the testing split. This data split involved leaving three drivers out for testing. For instance, ‘p002p014p012’ signifies that drivers with driverIDs ‘p002’, ‘p014’, and ‘p012’ were considered for the test set in the first fold. The remaining drivers were used for the training set. We follow this procedure eight times, each time keeping different groups of drivers separate for testing. We conduct 8-fold driver-level cross-validation on the StateFarm dataset. The other columns display the Top-1 accuracy of each model for different test splits. The last row provides an average Top-1 accuracy for each model across all folds. Amongst all the models, the Single-frameCLIP model with ViT-L/14 architecture emerged as the best performing, achieving an average Top-1 accuracy of 83.15%. Fig. 6 visually represents these results. The bars illustrate the average Top-1 accuracy of the models, while the vertical lines on the top indicate the variance in Top-1 accuracies across all eight folds.
TABLE VII: Single-frameCLIP vs. Traditional models:Top-1 accuracy on StateFarm dataset
Test DriverID ResNet18 MobileNet-v2 AlexNet ViT-L/14
p002p014p012 72.81 74.09 68.23 75.66
p015p016p021 81.35 82.16 75.33 88.12
p024p022p026 87.11 86.16 81.67 89.03
p041p039p035 87.17 85.60 87.12 90.27
p042p045p047 84.00 81.91 81.72 89.03
p051p049p050 88.72 81.48 70.38 78.13
p061p056p052 86.17 82.50 75.16 72.39
p066p072p064 74.14 66.45 59.14 71.46
Avg. Top-1 82.68 80.04 74.84 83.15
Figure 6: The average Top-1 accuracy of Single-frameCLIP and other baselines on the StateFarm dataset. We perform 8-fold cross-validation and show the mean and variance of Top-1 accuracy values across all the folds along the Y-axis.
TABLE VIII: Single-frameCLIP vs. traditional modelsTop-1 accuracy on SynDD1 dataset
Average Leaving
Models Backbones Top-1 a set of
accuracy driver out
Single-frameCLIP ViT-L/14 57.347
ViT-B/16 48.940
RN101 48.521
ViT-B/32 35.382
Traditional Models ResNet18 42.621
MobileNet-v2 36.102
AlexNet 13.712
The results obtained with the eight distracted driving classes in the SynDD1 dataset are presented in Table VIII. Similar to the experiments on the StateFarm dataset, this experiment involved 7-fold driver-level cross-validation with ‘leaving two drivers out’ for testing in each fold. The third column in Table VIII represents the average Top-1 accuracy across all the 7-folds. The results reveal that the ViT-L/14 model outperformed others, achieving a Top-1 accuracy of 57.347% (average accuracy across 7-folds). The remaining models followed with lower accuracies, with ViT-B/16 achieving 48.940% and RN101 at 48.521%. In conclusion, the Single-frameCLIP model with the ViT-L/14 architecture demonstrated superior performance in terms of Top-1 accuracy both in the StateFarm and SynDD1 datasets.
@@
In an effort to analyze the performance of the Single-frame CLIP model with smaller amounts of training data, a series of experiments were conducted utilizing a progressively reduced training dataset. In this experiment, we train the traditional models and our proposed CLIP-based frameworks in a similar experimental setting on the StateFarm dataset. We separate 3 drivers (driverID p012, p014, p021) for testing and progressively removed four drivers from the training set to create a smaller training set, which is $80\%$, $60\%$, $40\%$, and $20\%$ of the original training set. For this process, instead of randomly removing frames, we reduce the training data based on the driverID since, in real-world scenarios, the model will be expected to generalize its learning to new, unseen drivers. The percentage of training data, the number of drivers present in those splits, and the number of samples are shown in Table  IX. For this experiment on the StateFarm dataset, we employ an 8-fold cross-validation scheme where we excluded three drivers (driverID p012, p014, p021) from every fold and used them as test subjects. We picked the ViT/L-14 backbone for the Single-frameCLIP. All the models used the same training samples and were finally tested on the same test set. The outcomes are presented in Fig. 7. The results suggest that reducing the size of the training set does not necessarily compromise the Single-frameCLIP model’s performance. Unlike the other models assessed, which displayed a decrease in accuracy, the Single-frameCLIP model maintained its performance levels consistently even when the quantity of training data was reduced to 20% of the original amount.
TABLE IX: Training Datasize for reduced training set
@@
Figure 7: Performance comparison of the Single-frameCLIP model and traditional models trained on varying proportions of the StateFarm dataset. Training sets of 80%, 60%, 40%, and 20% were created by progressively removing drivers based on their driverID. The figure reports the models’ average accuracies obtained through an 8-fold cross-validation scheme based on driverID.
@@
TABLE X: Multi-frameCLIP vs. traditional models(at Sampling rate 1FPS)
@@
We provide an extensive performance comparison of Multi-frameCLIP with traditional models on three different datasets in Table XIII. To ensure consistency and reliability in our experiments, we run the listed models at a 1FPS sampling rate with the same 7-fold cross-validation and ‘leaving a set of driver out’ settings across all the datasets. The Multi-frameCLIP model shows superior performance compared to the other baselines, achieving a Top-1 accuracy of 93.39% on DMD, 89.14% on SAM-DD, and 70.36% on the SynDD1 dataset. These results are significantly higher than those of models such as AlexNet, VGG16, MobileNet-v2, and ResNet variants when tested under the same conditions. In addition, it showcases the Multi-frameCLIP model’s efficacy in capturing driver behavior in different scenes, making it a suitable benchmark for distracted driving detection applications. However, it is important to note that accuracy fluctuates with changes in hyperparameters such as model type, sampling rate, camera view, and so on (discussed in the following sections), indicating that these are significant factors to consider to get an optimized performance.
@@
We showcase a comparative analysis of the VideoCLIP against the traditional baselines in Table XI. It shows that VideoCLIP gets significant performance gain and outperforms its counterparts across three different datasets. On the DMD dataset, VideoCLIP outperforms all other models, including advanced configurations like MobileNet-v2 (MLP), 3D CNN variants and modalities, achieving a Top-1 accuracy of 98.44%. Similarly, in the SAM-DD dataset, VideoCLIP tops the chart with 97.86% accuracy, surpassing other SOTA models such as Swin Transformers and ViTs. Furthermore, on the SynDD1 dataset, VideoCLIP maintains its lead with an 81.85% accuracy at a 20FPS rate, which is substantially higher than the average Top-1 accuracy of the competing model [69]. These results highlight the efficacy of the VideoCLIP model in harnessing RGB modality and underscore its robustness across different datasets within the ’leave a set of driver out’ setting.
TABLE XI: VideoCLIP vs. traditional models(with Sampling rate 1FPS)
@@
*best performing model proposed by the author
**average Top-1 accuracy for the same classes used in VideoCLIP
TABLE XII: Single-frameCLIP: F1-score on DMD and SAM-DD
@@
TABLE XIII: Multi-frameCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)

### 21 Mar 2024 v4

@@ V Results
V-A Performance summary of the CLIP-based Frameworks
TABLE III: Performance Summary of CLIP-based
@@ V-A Performance summary of the CLIP-based Frameworks
(ours) Top-1 Top-1 Top-1 Top-1
Zero-shotCLIP 54.00 58.22 48.12 66.52
Single-frameCLIP 57.35 83.15 82.65 89.14
Multi-frameCLIP 70.36 N/A 93.39 89.14
VideoCLIP 81.85 N/A 98.44 97.86
We present the performance summary of various CLIP-based frameworks on the four distracted driving datasets in Table III. The VideoCLIP model with an optimal hyperparameter setting demonstrated the most superior performance with a Top-1 accuracy of 81.85%, 98.44% and 97.86% on SynDD1, DMD and SAM-DD datasets, respectively. The multi-frameCLIP framework also showed notable results with its optimal hyperparameter setting. Single-frameCLIP and Zero-shotCLIP reported lower Top-1 accuracies on the SynDD1 dataset compared to the SAM-DD dataset due to the dataset size. On the StateFarm dataset, only Single-frameCLIP and Zero-shotCLIP were evaluated due to the absence of video data, which is a prerequisite for Multi-frameCLIP and VideoCLIP. Single-frameCLIP outperformed Zero-shotCLIP with optimal hyperparameter setting, achieving a Top-1 accuracy of 83.15%. These results highlight the effectiveness of VideoCLIP and Multi-frameCLIP in leveraging temporal information from video data for improved performance.
V-B Performance of the Frameworks with Optimal Hyperparameter Setting
TABLE IV: Proposed frameworks performance with Optimal Hyperparameter setting
Framework Dataset CLIP Camera FPS Top-1 Leaving a set
name backbone view accuracy of driver out
Zero-shotCLIP DMD ViT-L/14@336px Side-view N/A 48.12 N/A
SAM-DD ViT-L/14@336px Side-view N/A 66.52 N/A
StateFarm ViT-L/14@336px Side-view N/A 58.22 N/A
SynDD1 ViT-L/14@336px Dashboard N/A 54.00 N/A
Single-frameCLIP DMD ViT-L/14@336px Side-view N/A 82.65
SAM-DD ViT-L/14 Side-view N/A 89.14
StateFarm ViT-L/14 Side-view N/A 83.15
SynDD1 ViT-L/14 Dashboard N/A 57.35
Multi-frameCLIP DMD ViT-L/14@336px Side-view 1 93.39
SAM-DD ViT-L/14 Side-view 1 89.14
SynDD1 ViT-L/14 Dashboard 1 70.36
VideoCLIP DMD ViT-L/14@336px Side-view 1 98.44
SAM-DD ViT-L/14 Side-view 1 97.86
SynDD1 ViT-L/14 Dashboard 20 81.85
Table IV summarizes the performance of Zero-shotCLIP, Single-frameCLIP, Multi-frameCLIP and VideoCLIP—across various datasets with optimal hyperparameter settings. Each framework achieves its best results using either the ViT-L/14 or the ViT-L/14@336px backbone, evaluated on specific datasets. Performance is measured in terms of Top-1 accuracy, with all experiments incorporating a “leaving a set of driver out” and a $k$-fold cross-validation step. Zero-shotCLIP serves as a benchmark for initial model results. Single-frameCLIP demonstrates consistent performance with accuracies ranging from 57.35% to 89.14% across different views and datasets. Multi-frameCLIP exhibits performance enhancements, highlighting a significant accuracy increase to 93.39% on the DMD dataset with a side-view camera setting. VideoCLIP surpasses the other models, achieving peak accuracy scores on all three datasets. It demonstrates the significant advantage of leveraging temporal features and video-based feature representation. This result highlights the effectiveness of adopting the VideoCLIP model and the substantial improvements possible by incorporating video-based features over frame-based methods. It also illustrates how the selection of optimal hyperparameters such as camera views, sampling rates, and CLIP backbones, plays a pivotal role in maximizing model performance.
V-C Performance Comparison with Other Works
@@
In this section, we compare the performance of the Single-frameCLIP model with the traditional distracted driving models on frame-based datasets.
TABLE V: Single-frameCLIP vs. traditional models
Dataset Models Top-1 Leaving a set
accuracy of driver out
StateFarm Single-frameCLIP 83.15
ResNet18 82.68
MobileNet-v2 80.04
AlexNet 74.84
SynDD1 Single-frameCLIP 57.35
ResNet18 42.62
MobileNet-v2 36.10
AlexNet 13.71
Table V shows the performance of the Single-frameCLIP model against traditional Convolutional Neural Network-based (CNN) baselines on two frame-based datasets (StateFarm and SynDD1) in terms of Top-1 accuracy. The Single-frameCLIP based on ViT-L/14 backbone outperforms the traditional CNN models on both datasets. These results suggest that the Single-frameCLIP is more effective at leveraging contextual information and understanding complex visual patterns compared to traditional CNNs. It also demonstrates that the Single-frameCLIP generalizes better to unseen drivers (as indicated by the “leaving a set of driver out” column) and diverse driving scenarios, making it a more effective framework for analyzing driver behavior. Moreover, Fig. 3 represents the mean and variance of Top-1 accuracy of the Single-frameCLIP and CNN baseline models on the StateFarm dataset. The bars illustrate the average Top-1 accuracy of the models, while the vertical lines on the top indicate the variance in Top-1 accuracies across all the folds. This barplot highlights the consistent performance of the Single-frameCLIP model in distracted driving detection.
Figure 3: The mean and variance of Top-1 accuracy of the Single-frameCLIP and CNN baselines on the StateFarm dataset.
@@
In an effort to analyze the performance of the Single-frame CLIP model with smaller amounts of training data, a series of experiments are conducted utilizing progressively reduced training datasets. In this experiment, we train the traditional CNN models and our proposed Single-frameCLIP model with a smaller training data sampled from the StateFarm dataset. We separate 3 drivers (driverID p012, p014, p021) for testing and progressively removed four drivers from the training set to create a smaller training set, which is $80\%$, $60\%$, $40\%$, and $20\%$ of the original training set. In this procedure, rather than indiscriminately eliminating random frames, we selectively remove the training samples according to driverID. This approach mirrors real-world scenarios where the model must adapt its learned behaviors to new and unseen drivers. The reduced training data, number of drivers present in the training splits, and the number of samples are shown in Table VI.
TABLE VI: Training Data for the reduced training set Experiment
@@
Figure 4: Performance comparison of the Single-frameCLIP model and traditional CNN models trained on varying proportions of the StateFarm dataset. It shows average accuracies for 8-fold cross-validation.
For this experiment, we employ 8-fold cross-validation and excluded three drivers (driverIDs are on the rightmost column) from every fold and used them as test subjects. We utilize the CLIP ViT/L-14 as the backbone of Single-frameCLIP. All the models use the same training samples and are finally tested on the same test set. The results in Fig. 4 suggest that reducing the size of the training set does not necessarily compromise the Single-frameCLIP model’s performance. Unlike the other models, which displayed a decrease in accuracy, the Single-frameCLIP model maintained its performance levels consistently even when the quantity of training data was reduced to 20% of the original size.
@@
TABLE VII: Multi-frameCLIP vs. traditional models(at Sampling rate 1FPS)
@@
We provide an extensive performance comparison of Multi-frameCLIP with traditional models on three video datasets in Table VII. To ensure consistency and reliability in our comparison, we run the listed models at a 1FPS sampling rate with the same 7-fold cross-validation and “leaving a set of driver out” settings across all the datasets. The Multi-frameCLIP model shows superior performance compared to the recent baseline models, achieving a Top-1 accuracy of 93.39% on DMD, 89.14% on SAM-DD, and 70.36% on the SynDD1 dataset. These results are significantly higher than the recent baselines when tested under the same setting. In addition, it showcases the Multi-frameCLIP model’s efficacy in capturing driver behavior in different scenes, making it a suitable benchmark. However, it is important to note that accuracy fluctuates with changes in hyperparameters such as model type, sampling rate, camera view, and so on, indicating that these are significant factors to consider to get an optimized performance.
@@
TABLE VIII: VideoCLIP vs. traditional models(at Sampling rate 1FPS)
@@
*average Top-1 accuracy for the same classes used in VideoCLIP
Table VIII showcases a comparative analysis of the VideoCLIP against the recent baseline models. It shows that VideoCLIP gets significant performance gain and outperforms its counterparts across three different datasets. For example, in the DMD dataset, VideoCLIP outperforms the other models, including MobileNet-v2 (MLP), and 3D CNN variants, achieving a Top-1 accuracy of 98.44%. Similarly, in the SAM-DD dataset, VideoCLIP tops the chart with 97.86% accuracy, surpassing other state-of-the-art models such as ViT and VGG16. These results highlight the efficacy of the VideoCLIP model in harnessing the RGB modality of the input video and underscore its robustness across different video datasets.
TABLE IX: Single-frameCLIP: F1-score on DMD and SAM-DD
@@
V-D Hyperparameters and Ablation Studies
1)  Class-level Analysis
We illustrate the class-level F1-scores of Single-frameCLIP on the DMD and SAM-DD datasets in Table IX to understand the most challenging actions to classify. The average F1-scores on DMD and SAM-DD datasets are 0.819 and 0.847, respectively. In addition, we demonstrate the confusion matrix of Single-frameCLIP on fold02 of the DMD dataset as an example in Fig. 5. We have found that the DMD classes such as “reaching to backseat”, “reaching to side” and “driving safely” exhibited lower F1-scores ($<$0.70), indicating that these categories pose greater challenges for the framework to classify correctly. These action classes appeared to generate similar image features, leading to a higher rate of misclassification. Additionally, we noticed that only two SAM-DD classes “head dropping” and “touching hair” got F1-scores below 0.70 among the ten classes. To further illustrate this, Fig. 6 provides visual examples of instances where the Single-frameCLIP model incorrectly classified actions, comparing the model’s outputs against the ground truth.
TABLE X: Single-frame CLIP: Removing vs. merging classes results on SynDD1 dataset
Removing 5 3 0.64 0.71 0.84 0.82 0.80 - 0.763 $75.62\%$
Merging 6 3 0.51 0.50 0.82 0.75 0.88 0.73 0.698 $69.41\%$
Removing with
dashboard-view
5 1 0.77 0.87 0.89 0.94 0.89 - 0.873 87.08%
TABLE XI: Multi-frameCLIP at different sampling rates on SynDD1 dataset (with leaving a set of driver out)